# 3D-MolT5: Towards Unified 3D Molecule-Text Modeling with 3D Molecular Tokenization

### v1

@@ Appendix B Pre-training
The pre-training is done on eight NVIDIA 80GB A100 GPUs. The total number of steps for pre-training is 400,000, with warm-up steps set to 10,000. AdamW [31] with Root Mean Square (RMC) scaling optimizer is used. The peak learning rate is 2e-3 with cosine decay, and the minimum learning rate is 1e-5. The maximum length for input and output is 512. The batch size is set to 768. As shown in Table 6, the sizes of pre-training datasets vary significantly. To balance the data from different tasks during pre-training, we implement a batch-level balancing strategy. Each batch evenly includes data from all tasks, ensuring a more balanced and comprehensive pre-training process. For smaller datasets, such as molecule-text pairs from PubChem, we employ a round-robin strategy to repeat their usage multiple times, compensating for their limited size. For all molecular data, we first get its canonical SMILES from the provided SMILES or 3D structure using RDKit [22], and then convert it to SELFIES using selfies toolkit [^5] https://github.com/aspuru-guzik-group/selfies [20]. The resulting SELFIES are also wrapped by special tokens $\langle bom\rangle$ and $\langle eom\rangle$ to differentiate from text.
@@
Table 6: Statistics of pre-training datasets. Deno. refers to T5 [43] denoising task; Tran. refers to translation task. For the PubChem dataset, the 3D structure is obtained using the MMFF algorithm in RDKit [22] and the text enriched by GPT-3.5 [36].
@@
Appendix C Fine-tuning
@@
Details for datasets for fine-tuning are shown in Table 7. For all downstream datasets, we follow the same pipeline as described in Appendix B to first get the canonical SMILES from the provided SMILES or 3D structure using RDKit [22], and then convert it to SELFIES wrapped by $\langle bom\rangle$ and $\langle eom\rangle$. All reported results for 3D-MolT5 are the mean value obtained from three independent random runs.
Table 7: Dataset statistics for donwstream fine-tuning. All the datasets are in instruction format. Small differences exist between our processed datasets and the original version, as we discard the data that can not be processed by E3FP [2].
@@
For PubChemQC [34] and PubChem [19] datasets, we use the instruction versions built by 3D-MoLM [25]. The Generalist version of 3D-MolT5 here is trained simultaneously on these two datasets with three types of tasks: computed property prediction, description property prediction, and 3D molecule captioning. For descriptive property prediction, the descriptive text is generated by employing GPT-3.5 [36] to read molecular captions and create five QA pairs for each molecule. The reported baseline results are derived from 3D-MoLM [25]. Specifically, the baseline method 2D-MoLM is a variant of 3D-MoLM [25], where the 3D molecular encoder is replaced with a 2D molecular encoder. The baseline Llama2-7B [52] directly removes the 3D molecular encoder of 3D-MoLM [25] and uses 1D SMILES as the molecular representation.
For QM9 [47] dataset, we use its instruction version built by Mol-Instructions [14]. The 3D structures are downloaded from DeepChem [44]. The Generalist version for QM9 is trained on the combination of its three subsets: HOMO, LUMO, and HOMO-LUMO gap. The reported baseline results are derived from BioT5+ [40].
For CheBI-20 [13] dataset, we manually convert it to instruction version. To avoid data leakage, we exclude the molecules of the CheBI-20 test set that are also present in the PubChem 3D molecule-text pairs in pre-training. In this task, molecular names are removed from the text to prevent the model from learning a simple mapping from molecular names to 1D sequences. The reported baseline results are mainly sourced from MolT5 [13], MolReGPT [23], MolFM [33], GIT-Mol [27], MolXPT [29], and BioT5 [39].
Appendix D Case Study
@@
Appendix E Limitations
In 3D-MolT5, the 3D structure information is only incorporated in the input, and 3D-MolT5 can not generate 3D molecular structure directly. The molecule-text pairs for pre-training are in limited size (less than 300K). Larger datasets with high-quality molecule-text pairs may further enhance the performance of 3D-MolT5. We leave these constraints for future exploration.

### v2

@@ E.2 Data and Configuration
The pre-training is done on eight NVIDIA 80GB A100 GPUs. The total number of steps for pre-training is 400,000, with warm-up steps set to 10,000. AdamW (Loshchilov & Hutter, 2019) with Root Mean Square (RMC) scaling optimizer is used. The peak learning rate is 2e-3 with cosine decay, and the minimum learning rate is 1e-5. The maximum length for input and output is 512. The batch size is set to 768. As shown in Table 10, the sizes of pre-training datasets vary significantly. To balance the data from different tasks during pre-training, we implement a batch-level balancing strategy. Each batch evenly includes data from all tasks, ensuring a more balanced and comprehensive pre-training process. For smaller datasets, such as molecule-text pairs from PubChem, we employ a round-robin strategy to repeat their usage multiple times, compensating for their limited size. For all molecular data, we first get its canonical SMILES from the provided SMILES or 3D structure using RDKit (Landrum et al., 2023), and then convert it to SELFIES using selfies toolkit (Krenn et al., 2020). [^6] https://github.com/aspuru-guzik-group/selfies The resulting SELFIES are also wrapped by special tokens $\langle bom\rangle$ and $\langle eom\rangle$ to differentiate from text.
@@
Table 10: Statistics of pre-training datasets. Deno. refers to T5 (Raffel et al., 2020) denoising task; Tran. refers to translation task. For the PubChem dataset, the 3D structure is obtained using the MMFF algorithm in RDKit (Landrum et al., 2023) and the text enriched by GPT-3.5 (OpenAI, 2023).
@@
Appendix F Fine-tuning
@@
Details for datasets for fine-tuning are shown in Table 11. For all downstream datasets, we follow the same pipeline as described in Appendix E to first get the canonical SMILES from the provided SMILES or 3D structure using RDKit (Landrum et al., 2023), and then convert it to SELFIES wrapped by $\langle bom\rangle$ and $\langle eom\rangle$. All reported results for 3D-MolT5 are the mean value obtained from three independent random runs.
Table 11: Dataset statistics for donwstream fine-tuning. All the datasets are in instruction format. Small differences exist between our processed datasets and the original version, as we discard the data that can not be processed by E3FP (Axen et al., 2017).
@@
For PubChemQC (Maho, 2015) and PubChem (Kim et al., 2019) datasets, we use the instruction versions built by 3D-MoLM (Li et al., 2023c). The Generalist version of 3D-MolT5 here is trained simultaneously on these two datasets with three types of tasks: computed property prediction, description property prediction, and 3D molecule captioning. We follow the same sampling algorithm as 3D-MoLM (Li et al., 2023c), where the sampling probabilities for each task are proportional to the fourth root of the size of its data. For descriptive property prediction, the descriptive text is generated by employing GPT-3.5 (OpenAI, 2023) to read molecular captions and create five QA pairs for each molecule. The reported baseline results are derived from 3D-MoLM (Li et al., 2023c). Specifically, the baseline method 2D-MoLM is a variant of 3D-MoLM (Li et al., 2023c), where the 3D molecular encoder is replaced with a 2D molecular encoder. The baseline Llama2-7B (Touvron et al., 2023) directly removes the 3D molecular encoder of 3D-MoLM (Li et al., 2023c) and uses 1D SMILES as the molecular representation.
For QM9 (Ruddigkeit et al., 2012) dataset, we use its instruction version built by Mol-Instructions (Fang et al., 2023). The 3D structures are downloaded from DeepChem (Ramsundar et al., 2019). The Generalist version for QM9 is trained on the direct combination of its three subsets: HOMO, LUMO, and HOMO-LUMO gap. The reported baseline results are derived from BioT5+ (Pei et al., 2024a).
For CheBI-20 (Edwards et al., 2022) dataset, we manually convert it to instruction version. To avoid data leakage, we exclude the molecules of the CheBI-20 test set that are also present in the PubChem 3D molecule-text pairs in pre-training. In this task, molecular names are removed from the text to prevent the model from learning a simple mapping from molecular names to 1D sequences. The reported baseline results are mainly sourced from MolT5 (Edwards et al., 2022), MolReGPT (Li et al., 2023a), MolFM (Luo et al., 2023), GIT-Mol (Liu et al., 2024), MolXPT (Liu et al., 2023b), and BioT5 (Pei et al., 2023).
Appendix G Case Study
@@
Appendix H Limitations
In 3D-MolT5, the 3D structure information is only incorporated in the input, and 3D-MolT5 can not generate 3D molecular structure directly, which is mainly caused by two factors. (1) The hash algorithm and “folding” process are irreversible and may introduce value collisions, though the probability is small. (2) The pooling of embeddings at each iteration into the 3D embedding, and the pooling of 1D and 3D embeddings. The unified modeling of both the understanding and generation of 3D molecular structures remains an area for future exploration.